# ELMS: Elasticized Large Language Models On Mobile Devices

### v1

@@ Abstract.
On-device Large Language Models (LLMs) are revolutionizing mobile AI, enabling applications such as UI automation while addressing privacy concerns. Currently, the standard approach involves deploying a single, robust LLM as a universal solution for various applications, often referred to as LLM-as-a-Service (LLMaaS). However, this approach faces a significant system challenge: existing LLMs lack the flexibility to accommodate the diverse Service-Level Objectives (SLOs) regarding inference latency across different applications. To address this issue, we introduce ELMS, an on-device LLM service designed to provide elasticity in both the model and prompt dimensions of an LLMaaS. This system includes: A one-time neuron reordering technique, which utilizes the inherent permutation consistency within transformer models to create high-quality, elastic sub-models with minimal runtime switching costs. A dual-head compact language model, which efficiently refines prompts and coordinates the elastic adaptation between the model and the prompt. We have implemented this elastic on-device LLM service on several off-the-shelf (COTS) smartphones and evaluate ELMS using both standalone NLP/mobile-agent datasets and synthesized end-to-end traces. Across a range of SLOs, ELMS surpasses four strong baselines by up to 16.83% and 11.04% in absolute accuracy on average, with less than 1% Time-To-First-Token (TTFT) switching overhead, comparable memory usage, and fewer than 100 offline GPU hours.
copyright: none
@@ 1. Introduction
Large Language Models (LLMs) are ushering in a transformative era for mobile AI. A multitude of killer apps are built on top of LLMs, encompassing mobile UI automation (Wen et al., 2024), API-calling (Chen and Li, 2024) and screen content comprehension (rew, 2024). For instance, one can easily place an order by simply saying “Order a pizza now from Pizza Hut” in smartphone.
With ever-growing privacy concerns, there is an emerging paradigm of mobile LLM deployment: local LLM as a system service, namely LLM-as-a-Service (LLMaaS) (Yin et al., 2024; Li et al., 2024; AIC, 2024; Mei et al., 2024; Packer et al., 2024). Akin to location or notification services, LLM service runs as a standalone system service in background and serves LLM requests (text in, text out) from third-party apps. Such a paradigm is completely feasible thanks to the world knowledge and in-context learning abilities of LLMs. With the OS’s visibility into the service, LLMaaS is also resource-efficient. Only one copy of LLM weights is needed in memory, preventing the device memory from being exhausted by app-specific models. With a converged model architecture and operator set, the LLM service can better enjoy the system level scheduling optimizations (e.g., batching or priority queues) and hardware accelerators (e.g., GPUs, DSPs, or NPUs). An industrial example is Google’s Android AICore (AIC, 2024), a built-in on-device LLM service in Android OS that has been used by apps like GBoard smart reply and Pixel voice recorder. Similar paradigm can also be found in Apple (App, 2024) and Intel (int, 2024).
@@
Elastic on-device LLM service. We identify a crucial yet unexplored system feature of on-device LLM service — performance elasticity. Specifically, each app calling LLM service demands its own Service-Level Objective (SLO) on inference latency, yet a single static LLM cannot meet the diversified SLOs. Demand of elasticity is further exaggerated and complicated since LLM inference consists of two distinct stages: prefill (prompt processing speed) and decode (token generation speed), whose latencies are measured by Time-To-First-Token (TTFT) and Time-Per-Output-Token (TPOT), respectively. For instance, a chatbot (lla, 2024) must behave both low TTFT and low TPOT to match human reading speed; a UI-automation agent (Zhang et al., 2024; Wen et al., 2024) typically requires a low TTFT and an acceptable TPOT, as TPOT can be overlapped with UI manipulations; a screen-event recorder (rew, 2024) running in background only needs a tolerable TTFT/TPOT. Failing to meet an SLO leads to serious consequences: a significant degradation of user experience, or failure in the interactions between LLM agents and the environments/tools (Zhang et al., 2024; Chen and Li, 2024; App, 2024).
@@
In a realistic LLM service with elastic feature, an app sends a prompt to the service as well as the SLO expected, and the LLM service needs to provide the highest text generation quality without failing the SLO. Fundamentally differing from prior work that elasticize CNNs in pre-LLM era (Wen et al., 2023; Han et al., 2021; Fang et al., 2018, 2023), a key opportunity in our system model is that, both model and input (prompts) dimensions of the LLM can be elasticized. Through multi-scale LLM pruning technique (Xia et al., 2024; Ma et al., 2023; Sun et al., 2024), one can get a crucial subset of weights that running at different speed; similarly, through a small language model that scores the importance of each input token (Jiang et al., 2024; Pan et al., 2024; Modarressi et al., 2022), one can prune the prompt into different lengths on demand. In both ways, LLM accuracy is sacrificed for faster text generation in a flexible manner. Our pilot experiments in §$\S$2.2 explore how each dimension of elasticity impacts the LLM inference latency: TTFT (often more time consuming and requires higher elasticity) is proportional to both the prompt length and the model size; TPOT is mainly proportional to the model size with the help of KV cache (Kwon et al., 2023).
Challenges. However, designing an LLMaaS with efficient elasticity faces the following unique challenges.
@@
$\bullet$ Sensitive prompt-model orchestration strategy. There exist multiple elastic strategies to meet an SLO, yet their output text quality could differ significantly. Exemplified with a real prompt from ARC_E dataset in Figure 5, although both elastic strategies (50%/20% prompt/model pruning vs. 20%/50% prompt model pruning) can meet the SLO, only the first strategy leads to a correct generated answer. Another instance is that with randomized strategy, the top5 API selection of Octopus dataset exhibits a 15.2% accuracy loss to the oracle strategy on an SLO with 50% TTFT and 80% TPOT of the full LLaMA-7B model. How to orchestrate the two dimensions of elastification to maximize the LLM output quality at request level has not been touched in prior study.
@@
Our solution: ELMS. We present ELMS, a first-of-its-kind elastic on-device LLM service that tackles the above challenges through following novel techniques.
@@
One-shot reordering of permutation consistent units (§$\S$3.2). This technique performs pruning on a fine granularity of permutation consistent units in Transformers. Identified by ELMS, these units can be offline arbitrarily layouted in a block (e.g., Attention or MLP) while guarantee the equivalent input/output as original block, thereby fundamentally avoiding the runtime switching overhead. Specifically, such a unit is an entire attention head (i.e., columns in $W_{Q}/W_{K}/W_{V}$ and rows in $W_{O}$ with the same indices), or an entire MLP neuron (i.e., a columns in $W_{up}$ and a row in $W_{down}$). In the one-shot reordering, ELMS first profiles the importance of the units through an eXplainable-AI (XAI) (Huang et al., 2023; Selvaraju et al., 2019; Sundararajan et al., 2017) guided method, which measures unit importance via gradients. Then, ELMS reorders the units in memory based on their importance, making each contiguous memory segment (starting from base address to a memory pointer) a sub-model. After reordering by importance, the pruning always extends from the periphery inward, thus requiring no online reordering. The profiling and reordering are all done offline, incurring no online overhead. The online switching of sub-models is zero-cost by only moving the pointer. ELMS further incorporates other optimizations such as LoRA recovery and anchor layers locking to improve the sub-models quality.
@@
Dual-head Tiny Language Model (TLM) for prompt-model elastification orchestration (§$\S$3.3). Different from prior work (Pan et al., 2024; Jiang et al., 2024) that only focus on identifying important tokens, ELMS designs a dual-head TLM as an end-to-end solution for determining an optimal prompt-model elastification strategy. Specifically, the TLM features two heads, namely score-head and decision-head. During inference, the score-head estimates the token importance, while the decision-head chooses the proper prompt- and model- elastification ratios (mapped to the real latency by a one-shot profiling on various SoCs). At offline, the decision-head learns from the prompts and groundtruth of a generic corpora. As a self-induced labelling process, all possible strategies of a prompt with its SLOs are traversed to get the inference result, and an optimal strategy is recorded as the label. Both the training and inference of TLM are cost-effective with careful optimizations like reusing the compact MobileBert and bottom layers sharing.
Evaluation. We have fully implemented ELMS on 3 COTS smartphones and 5 base/instruction-tuned LLMs with 3B–7B parameters. The datasets encompass fundamental language tasks ARC_E/OBQA (Clark et al., 2018; Mihaylov et al., 2018), mobile UI automation tasks LlamaTouch (Zhang et al., 2024), and mobile system API-calling tasks Octopus (Chen and Li, 2024). Evaluation shows that ELMS achieves up to 16.83% and on average 11.04% higher absolute accuracy on end-to-end traces, and up to 34% and on average 22% on standalone datasets, when meeting all request SLOs. Within 2% absolute accuracy loss, ELMS can speed up TTFT by up to 5$\times$ and TPOP by up to 2$\times$. The runtime memory consumption is also on-par with non-elastic LLM service. The entire elastification of ELMS only takes 68.3 GPU hours (within \$100 GPU renting cost (hfg, 2024)) when elastictizing LLaMA-7B for MI14 smartphone, being affordable for most LLMaaS developers.
@@
  • We present the first elastic LLM service, ELMS, which fully exploits the space of model and prompt elastification through two novel techniques/designs, one-shot reordering of permutation consistent units and dual-head tiny language model.
  • We conduct comprehensive experiments on ELMS that demonstrate its superior performance over competitive baselines. The code will be made public after acceptance.
@@ 2.1. Elastic on-device LLM service
Figure 1. System model of elastic LLM service. The service can be upgraded/downgraded at runtime to adapt to various SLOs.
On-device LLM as a service. With ever-growing privacy concerns, recently there is a trend proposing running a single yet powerful (e.g., 7B-size) LLM as a system service (LLMaaS) on mobile devices (Yin et al., 2024; Li et al., 2024; AIC, 2024; Mei et al., 2024; Packer et al., 2024). In doing so, various apps/agents share the same LLM, minimizing the efforts and resources of preparing and managing task-specific models.
We identify a key system challenge of LLMaaS: the LLM requests from different apps or tasks have diversified SLO demands on inference latency, yet current LLMs lack the elasticity to do so. For instance, as listed in Table 1, a chatbot must behave both low TTFT and TPOT in order to match human reading speed. A UI-automation agent requires a relatively low TTFT to generate the first action and an acceptable TPOT, since the following latency can be overlapped with the manipulation of UI elements and thus transparent to users. Besides, these mobile-agents typically only decode few tokens compared to the prompt length. Failing to provide satisfactory latency for a request leads to serious consequences: a significant degradation in user experience, or failure in the interactions between LLM agents and the environment/tools.
@@
How to satisfy the heterogeneous demands of different LLM requests, while not degrade the LLM output quality significantly? One plausible solution is to deploy a dedicated-sized LLM for each SLO [^1] In cloud datacenters, a tighter SLO can be achieved by scaling up hardware resources, e.g., number of GPUs/TPUs. However, hardware resource of mobile devices is limited and not scalable.. This is unfriendly (and even infeasible) to both the LLMaaS developers and users. On one hand, costly GPU resources are required for pretraining multiple LLMs; on the other hand, memory consumption rises dramatically in order to manage these LLMs — running counter to the motivation behind LLMaaS.
System model of elastic LLM service. Thereby, we propose a model of elastic LLM service. As illustrated in Figure 1, at system layer, there is one running LLM that serves requests from the application layer (e.g., apps/agents). Each request consists of a prompt (a text sequence as LLM input) and an SLO (inference latency constraint). This single LLM can rapidly upgrade/downgrade itself to a more bulky/swift one at runtime to adapt to a specific SLO.
@@ 2.2. Opportunities and challenges
We present several observations to show the opportunities and challenges of realizing elastic LLM service.
@@
Observation#2: LLMs are elasticizable; yet elastic LLM does not necessarily translate to elastic LLM service. DNNs are known to be elasticizable: they can provide various compute-accuracy tradeoffs by a subset of their parameters (known as pruning (Fang et al., 2023; Ma et al., 2023)). For instance, Sheared-LLaMA (Xia et al., 2024) demonstrates that pruning off 60% parameters of a 7B-size LLM can still retain 74% of original accuracy on ARC_E dataset (Clark et al., 2018). LLMPruner (Ma et al., 2023) further shows that with lightweight Parameter-Efficient Fine-Tuning (PEFT) methods, the accuracy loss incurred by pruning could be recovered. Since pruning (and PEFT) generates sub-models that share the same superset of parameters, there is no overhead of extra memory or pre-training that mentioned in §$\S$2.1.
Yet, the challenge is that, since the model upgrades/downgrades itself to adapt to various requests’ SLOs, switching between these sub-models is not overhead-free. As shown in Figure 3a, although sub-models share the same superset of parameters, they are no longer contiguous in memory. One may change the deeply optimized dense kernels of on-device NN libraries (e.g., MNN (Jiang et al., 2020) or mllm (mll, 2024)) to handcrafted sparse kernels. However, these kernels typically undergo degraded performance without fully exploiting the parallelism of processors (e.g., mobile GPUs/NPUs or multi-core CPUs). Another compromising method is to perform a movement of parameter data for each model switching, as shown in Figure 3b. Although the switching overhead is mitigated from iteration/operator level to request level, it is still non-negligible. For instance, movement of LLaMA-7B’s a $W_{Q}$ matrix (4096$\times$4096) takes 139 ms on Redmi K60 Champion smartphone in the worst case, and consequently the entire model suffers time overhead at seconds level.
@@
However, the challenge is the sensitive prompt-model orchestration. An intuitive example is that, if a request sets its SLO as 40% TTFT and 80% TPOT [^3] See our formal definition of SLO in §$\S$3.1., we cannot know which strategy is golden a priori — a 50% prompt with an 80% model? an 80% prompt with a 50% model? or others? [^4] According to formula 1, there are multiple combinations of prompt and model that can meet the exemplified SLO. Prompts with various content naturally require distinct and customized strategies. In Figure 4 and Figure 5, we demonstrate that a strategy without careful design (e.g., random) may lead to a significant degradation on accuracy. We use LLaMA-7B elasticized by our method (elaborated in §$\S$3.2) on Octopus dataset. The prompt is elasticized by LLMLingua2.
Implications. On-device LLM service exhibits opportunities and challenges of being elasticized through prompt and model. In response, we present ELMS, an end-to-end elastic LLM service that orchestrates these two dimensions to provide latency-centric service.
3. ELMS Design

### v2

# Elastic On-Device LLM Service

@@ Abstract.
On-device Large Language Models (LLMs) are transforming mobile AI, catalyzing applications like UI automation without privacy concerns. Nowadays the common practice is to deploy a single yet powerful LLM as a general task solver for multiple requests. We identify a key system challenge in this paradigm: current LLMs lack the elasticity to serve requests that have diversified Service-Level Objectives (SLOs) on inference latency. To tackle this, we present ElastiLM, an on-device LLM service that elasticizes both the model and the prompt dimension of a full LLM. It incorporates (1) a one-shot neuron-reordering method, which leverages the intrinsic permutation consistency in transformer models to generate high-quality elasticized sub-models with minimal runtime switching overhead; (2) a dual-head tiny language model, which efficiently and effectively refines the prompt and orchestrates the elastification between model and prompt. We implement such an elastic on-device LLM service on multiple COTS smartphones, and evaluate ElastiLM on both standalone NLP/mobile-agent datasets and end-to-end synthesized traces. On diverse SLOs, ElastiLM outperforms 7 strong baselines in (absolute) accuracy by up to 14.83% and 10.45% on average, with ¡1% TTFT switching overhead, on-par memory consumption and ¡100 offline GPU hours.
On-Device Large Language Models, Elastic Service, Service-Level Objectives
#Corresponding author.
journalyear: 2025copyright: acmlicensedconference: The 31st Annual International Conference on Mobile Computing and Networking; November 4–8, 2025; Hong Kong, Chinabooktitle: The 31st Annual International Conference on Mobile Computing and Networking (ACM MOBICOM ’25), November 4–8, 2025, Hong Kong, Chinadoi: 10.1145/3680207.3765259isbn: 979-8-4007-1129-9/2025/11ccs: Computer systems organization Embedded and cyber-physical systemsccs: Computing methodologies Natural language generation
@@ 1. Introduction
Large Language Models (LLMs) are ushering in a transformative era for mobile AI. A multitude of killer apps are built on top of LLMs, encompassing mobile UI automation (Wen et al., 2024), API-calling (Chen and Li, 2024; Xie et al., 2024), and screen content comprehension (rew, 2024). For instance, one can easily place an order by simply saying “Order a pizza now from Pizza Hut” in smartphone.
With ever-growing privacy concerns, deploying LLMs on local devices (Yin et al., 2024; AIC, 2024; Mei et al., 2024; Packer et al., 2024) is attracting attentions increasingly. For instance, mobile GUI agent (Li et al., 2024) handles users’ requests by accessing the device screen information, which may contain highly private photos or chat history. To this end, Google has developed Android AICore (AIC, 2024), a built-in on-device LLM in Android OS that has been used by apps like GBoard smart reply and Pixel voice recorder.
@@
On-device LLM needs elasticity. LLM can handle ubiquitous language tasks, at the cost of large parameter size, making it neither “necessary” nor “practical” to deploy separated LLMs on a device. Instead, on-device LLM is often shared across tasks and apps, e.g., LLM-as-a-Service in OS (AIC, 2024). However, a single static LLM cannot meet the diversified Service-Level Objective (SLO) demanded by different LLM tasks. Consequently, on-device LLM needs elasticity. Demand of elasticity is further exaggerated and complicated since LLM inference consists of two distinct stages: prefill (prompt processing speed) and decode (token generation speed), whose latencies are measured by Time-To-First-Token (TTFT) and Time-Per-Output-Token (TPOT), respectively. For instance, a chatbot (lla, 2024) must behave both low TTFT and low TPOT to match human reading speed; a UI-automation agent (Zhang et al., 2024; Wen et al., 2024) typically requires a low TTFT and an acceptable TPOT, as TPOT can be overlapped with UI manipulations; a screen-event recorder (rew, 2024) running in background only needs a tolerable TTFT/TPOT. Even in a single app, the SLOs maybe also diversified due to the resource/task variation. Failing to meet an SLO leads to serious consequences: a significant degradation of user experience, or failure in the interactions between LLM agents and the environments/tools (Zhang et al., 2024; Chen and Li, 2024; App, 2024).
@@
Ideally, a request is sent to the elastic LLM with its prompt as well as the SLO expected, and the LLM needs to provide the highest text generation quality without failing the SLO. Fundamentally differing from prior work that elasticize CNNs in pre-LLM era (Wen et al., 2023; Han et al., 2021; Fang et al., 2018, 2023), a key opportunity in our system model is that, both model and input (prompts) dimensions of the LLM can be elasticized. Through multi-scale LLM pruning technique (Xia et al., 2024; Ma et al., 2023; Sun et al., 2024), one can get a crucial subset of weights that running at different speed; similarly, through scoring the importance of each input token (Jiang et al., 2024; Pan et al., 2024; Modarressi et al., 2022), one can prune the prompt into different lengths on demand. In both ways, LLM accuracy is sacrificed for faster text generation in a flexible manner. Our pilot experiments in $\S$2.2 explore how each dimension of elasticity impacts the LLM inference latency: TTFT (often more time consuming and requires higher elasticity) is proportional to both the prompt length and the model size; TPOT is mainly proportional to the model size with the help of KV cache (Kwon et al., 2023).
Challenges. However, elasticizing an on-device LLM faces the following unique challenges.
@@
$\bullet$ Sensitive prompt-model orchestration strategy. There exist multiple elastic strategies to meet an SLO, yet their output text quality could differ significantly. Exemplified with a real prompt from ARC_E dataset in Figure 5, although both elastic strategies (50%/20% prompt/model pruning vs. 20%/50% prompt/model pruning) can meet the SLO, only the first strategy leads to a correct generated answer. Another instance is that with randomized strategy, the top5 API selection of Octopus dataset exhibits a 15.2% accuracy loss to the oracle strategy on an SLO with 50% TTFT and 80% TPOT of the full LLaMA-7B model. How to orchestrate the two dimensions of elastification to maximize the LLM output quality at request level has not been touched in prior study.
@@
Our solution: ElastiLM. We present ElastiLM, a system for elasticizing on-device LLMs. It tackles the above challenges through the following novel techniques.
@@
One-shot reordering of permutation consistent units ($\S$3.2). This technique performs pruning on a fine granularity of permutation consistent units in Transformers. Identified by ElastiLM, these units can be offline arbitrarily layouted in a block (e.g., Attention or MLP) while guarantee the equivalent input/output as original block, thereby fundamentally avoiding the runtime switching overhead. Specifically, such a unit is an entire attention head (i.e., columns in $W_{Q}/W_{K}/W_{V}$ and rows in $W_{O}$ with the same indices), or an entire MLP neuron (i.e., a column in $W_{up}$ and a row in $W_{down}$). In the one-shot reordering, ElastiLM first profiles the importance of the units through an eXplainable-AI (XAI) (Huang et al., 2023; Selvaraju et al., 2019; Sundararajan et al., 2017) guided method, which measures unit importance via gradients. Then, ElastiLM reorders the units in memory based on their importance, making each contiguous memory segment (starting from base address to a memory pointer) a sub-model. After reordering by importance, the pruning always extends from the periphery inward, thus requiring no online reordering. The profiling and reordering are all done offline, incurring no online overhead. The online switching of sub-models is zero-cost by only moving the pointer. ElastiLM further incorporates other optimizations such as LoRA recovery and anchor layers locking to improve the sub-models quality.
@@
Dual-head Tiny Language Model (TLM) for prompt-model elastification orchestration ($\S$3.3). Different from prior work (Pan et al., 2024; Jiang et al., 2024) that only focus on identifying important tokens, ElastiLM designs a dual-head TLM as an end-to-end solution for determining an optimal prompt-model elastification strategy. Specifically, the TLM features two heads, namely score-head and decision-head. During inference, the score-head estimates the token importance, while the decision-head chooses the proper prompt- and model- elastification ratios (mapped to the real latency by a one-shot profiling on various SoCs). At offline, the decision-head learns from the prompts and groundtruth of a generic corpora. As a self-induced labelling process, all possible strategies of a prompt with its SLOs are traversed to get the inference result, and an optimal strategy is recorded as the label. Both the training and inference of TLM are cost-effective with careful optimizations like reusing the compact MobileBert and bottom layers sharing.
[…]
@@
  • We present ElastiLM, a system that fully exploits the space of model and prompt elastification through two novel techniques/designs, one-shot reordering of permutation consistent units and dual-head tiny language model.
  • We conduct comprehensive experiments on ElastiLM that demonstrate its superior performance over competitive baselines.
@@ 2.1. Elastic on-device LLM service
Figure 1. The system model. The service can be upgraded/downgraded at runtime to adapt to various SLOs.
On-device LLM needs elasticity. As listed in Table 1, a chatbot must behave both low TTFT and TPOT in order to match human reading speed. A UI-automation agent requires a relatively low TTFT to generate the first action and an acceptable TPOT, since the following latency can be overlapped with the manipulation of UI elements and thus transparent to users. Besides, these mobile-agents typically only decode few tokens compared to the prompt length, making a lower TTFT more important. Failing to provide satisfactory latency for a request leads to serious consequences: a significant degradation in user experience, or failure in the interactions between LLM agents and the environment/tools.
@@
How to satisfy the heterogeneous demands of different LLM requests, while not degrade the LLM output quality significantly? One plausible solution is to deploy a dedicated-sized LLM for each SLO [^1] In cloud datacenters, a tighter SLO can be achieved by scaling up hardware resources, e.g., number of GPUs/TPUs. However, hardware resource of mobile devices is limited and not scalable.. This is unfriendly (and even infeasible) to both the LLM service developers and users. On one hand, costly GPU resources are required for pretraining multiple LLMs; on the other hand, memory consumption rises dramatically in order to manage these LLMs — running counter to the motivation behind a single LLM service.
The system model. Thereby, we propose our system model. As illustrated in Figure 1, at system layer, there is one running LLM that serves requests from the application layer (e.g., apps/agents). Each request consists of a prompt (a text sequence as LLM input) and an SLO (inference latency constraint). This single LLM can rapidly upgrade/downgrade itself to a more bulky/swift one at runtime to adapt to a specific SLO.
@@ 2.2. Opportunities and challenges
Observation#2: LLMs are elasticizable; yet elastic LLM does not necessarily translate to elastic LLM service. DNNs are known to be elasticizable: they can provide various compute-accuracy tradeoffs by a subset of their parameters (known as pruning (Fang et al., 2023; Ma et al., 2023)). For instance, Sheared-LLaMA (Xia et al., 2024) demonstrates that pruning off 60% parameters of a 7B-size LLM can still retain 74% of original accuracy on ARC_E dataset (Clark et al., 2018). LLMPruner (Ma et al., 2023) further shows that with lightweight Parameter-Efficient Fine-Tuning (PEFT) methods, the accuracy loss incurred by pruning could be recovered. Since pruning (and PEFT) generates sub-models that share the same superset of parameters, there is no overhead of extra memory or pre-training that mentioned in $\S$2.1.
Yet, the challenge is that, since the model upgrades/downgrades itself to adapt to various requests’ SLOs, switching between these sub-models is not overhead-free. As shown in Figure 3a, although sub-models share the same superset of parameters, they are no longer contiguous in memory. One may change the deeply optimized dense kernels of on-device NN libraries (e.g., MNN (Jiang et al., 2020) or mllm (mll, 2024)) to handcrafted sparse kernels. However, these kernels typically undergo degraded performance without fully exploiting the parallelism of processors (e.g., mobile GPUs/NPUs or multi-core CPUs). Another compromising method is to perform a movement of parameter data for each model switching, as shown in Figure 3b. Although the switching overhead is mitigated from iteration/operator level to request level, it is still non-negligible. For instance, movement of LLaMA-7B’s a $W_{Q}$ matrix ($\times4096$) takes 139 ms on Redmi K60 Champion smartphone in the worst case, and consequently the entire model suffers time overhead at seconds level.
@@
However, the challenge is the sensitive prompt-model orchestration. An intuitive example is that, if a request sets its SLO as 40% TTFT and 80% TPOT [^3] See our formal definition of SLO in $\S$3.1., we cannot know which strategy is golden a priori — a 50% prompt with an 80% model? an 80% prompt with a 50% model? or others? [^4] According to formula 1, there are multiple combinations of prompt and model that can meet the exemplified SLO. Prompts with various content naturally require distinct and customized strategies. In Figure 4 and Figure 5, we demonstrate that a strategy without careful design (e.g., random) may lead to a significant degradation on accuracy. We use LLaMA-7B elasticized by our method (elaborated in $\S$3.2) on Octopus dataset. The prompt is elasticized by LLMLingua2.
3. ElastiLM Design